OpenStudy (anonymous):

A bicycle wheel with radius 26" rotates through an arc that measures 80°. What is the length of the arc of the tire that touched the ground? = ? PI inches

OpenStudy (jamesj):

If the wheel did a complete rotation--360 degrees--the arc length would 2pi.r = 2.pi.26 = 52.pi. Hence, having moved through just x degrees, the arc length is \[ \frac{x}{360} . 2\pi r \]

OpenStudy (jamesj):

In this case x = 80 and you also know r = 26 inches. Now solve.

OpenStudy (anonymous):

so what i do now like whats the answer lol sorry im in a rush?

OpenStudy (jamesj):

C'mon. Don't be so lazy. If x = 80, and r = 26, then Arc length = (x/360) . 2pi.r = (80/360).2pi(26) = ....

OpenStudy (anonymous):

still dont get it not being lazy but in a huge retricerush lol

OpenStudy (hrhsbull):

80/360*2*26pi = 11.56pi